IlU, the on product, product, Q Ksp U, 9. Given the following information: ZnCO3 (s) n (aq) CO3 (aq) -10 1.5 x HCO3 (aq) CO3 (aq) H (aq) Ka 5.6 x 10-11 What is the equilibrium constant for the reaction, Znco (s) H+(aq) 2+ HCO3 (aq) (aq) a. 8.4 x 10 21 b. 3.7 x 10-1 c. 2.7 d. 1.2 x 20 e. 2.1 x 10 -10
Answer #1

ZnCO3 (s)   <—> Zn2+(aq) + CO32- (aq) k1 = 1.5*10^-10

HCO3-(aq)   <—> CO32- (aq) + H+(aq) k2 = 5.6*10^-11

reverse the 2nd reaction:

ZnCO3 (s)   <—> Zn2+(aq) + CO32- (aq) k1 = 1.5*10^-10

CO32- (aq) + H+(aq) <—>HCO3-(aq) k3 = 1/k2

k3= 1/(5.6*10^-11) = 1.79*10^10

now add both reaction:

ZnCO3 (s) + H+(aq) <—> Zn2+(aq) +HCO3-(aq)

this is the required reaction

kc = k1*k3

= (1.5*10^-10)*(1.79*10^10)

= 2.7

Answer: c